### v1

# TurkBench: A Benchmark for Evaluating Turkish Large Language Models

Abstract
With the recent surge in the development of large language models, the need for comprehensive and language-specific evaluation benchmarks has become critical. While significant progress has been made in evaluating English-language models, benchmarks for other languages, particularly those with unique linguistic characteristics such as Turkish, remain less developed. Our study introduces TurkBench, a comprehensive benchmark designed to assess the capabilities of generative large language models in the Turkish language. TurkBench involves 8,151 data samples across 21 distinct subtasks. These are organized under six main categories of evaluation: Knowledge, Language Understanding, Reasoning, Content Moderation, Turkish Grammar and Vocabulary, and Instruction Following. The diverse range of tasks and the culturally relevant data would provide researchers and developers with a valuable tool for evaluating their models and identifying areas for improvement. We further publish our benchmark for online submissions at https://huggingface.co/turkbench
1 Introduction
In order to quantify the capabilities of large language models (LLMs), the research and development community relies on standardized evaluation frameworks, which are known as benchmarks. General Language Understanding Evaluation (GLUE) Wang et al. (2018), SuperGLUE Wang et al. (2019), Holistic Evaluation of Language Models (HELM) Liang et al. (2022), and the Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2021) are significant efforts in this regard. They provide a standardized set of tasks and metrics to assess models on diverse capabilities, from language comprehension and commonsense reasoning to solving mathematical problems and generating code.
Despite the success of these evaluation frameworks, a significant gap exists in their linguistic and cultural diversity. The majority of such benchmarks are predominantly English-centric, which stems from the availability of high-quality digital text in English Lai et al. (2024). This creates a critical challenge for evaluating models in other languages Park et al. (2024); Nacar et al. (2025); Baucells et al. (2025); Magnini et al. (2025), as direct translation of English benchmarks often fails to capture the unique linguistic structures, cultural nuances, and specific knowledge inherent to other languages Umutlu et al. (2025). Turkish, with its agglutinative morphology and distinct syntactic rules, presents unique challenges that cannot be easily assessed by frameworks designed for Indo-European languages Hakkani-Tür et al. (2002); Oflazer (2014).
Although there are efforts to develop benchmarks for Turkish LLMs Safaya et al. (2022); Uzunoğlu and Şahin (2023); Yüksel et al. (2024); Er et al. (2025), these resources primarily focus on exam-style tasks or do not cover many capabilities that drive real-world use of Turkish LLMs, such as open-ended instruction following, safety and content moderation, or fine-grained grammar and vocabulary control. Moreover, such benchmarks are mostly designed by using existing publicly available datasets, and lack of public leaderboards serving automated evaluation. In this study, we present TurkBench, which is a comprehensive benchmark specifically designed by curating novel data for the Turkish language. TurkBench provides a robust, culturally-aware, and linguistically-sound online evaluation tool to accurately measure the true capabilities of LLMs in Turkish.
The primary contribution of this study is the creation of a large-scale and comprehensive benchmark including 8,151 data samples across 21 distinct subtasks. These are organized under six main categories of evaluation: Knowledge, Language Understanding, Reasoning, Content Moderation, Turkish Grammar and Vocabulary, and Instruction Following. The data is obtained and crafted from high-quality Turkish materials through strategic partnerships with prestigious national institutions and university departments. We do not adapt existing datasets from the literature or create synthetic data. This ensures the tasks are not only challenging but also deeply embedded in the Turkish educational and cultural context.
All data in this benchmark are validated by human experts to follow three main criteria Umutlu et al. (2025): Correctness, Language Grammar, and Cultural Sensitivity (We provide the details of quality validation in Appendix 9.1). Evaluation metrics are designed for each task with accuracy being the primary measure for the classification tasks. More complex tasks, such as summarization and bias detection, utilize the LLM-as-a-Judge method Zheng et al. (2023). For the Semantic Textual Similarity task, Pearson and Spearman correlation coefficients are employed.
For researchers and developers, TurkBench provides an important tool to diagnose the strengths and weaknesses of their models, guiding targeted improvements and supporting innovation within the Turkish AI ecosystem. By establishing a standardized evaluation framework, TurkBench enables more meaningful and direct comparisons between different Turkish LLMs, which would promote a competitive environment that accelerates progress. Furthermore, the focus on content moderation and safety-related tasks such as bias and toxicity detection will support the development of more responsible and reliable AI systems for Turkish-speaking users. We also provide an online leaderboard [^1] https://huggingface.co/turkbench to evaluate Turkish models in our benchmark.
2 Related Work
LLM Benchmarking
Benchmarking large language models has evolved from single-task accuracy tests to broad suites that probe knowledge, reasoning, robustness, and safety. Measuring Massive Multitask Language Understanding (MMLU) evaluates multitask performance over 57 subjects, showing that scaling improves broad domain knowledge but also revealing persistent weaknesses on reasoning-heavy questions Hendrycks et al. (2021). General Language Understanding Evaluation (GLUE) Wang et al. (2018) and SuperGLUE Wang et al. (2019) aim to evaluate English-centric language understanding by providing public training sets and private test sets that can be assessed through an evaluation server. BIG-Bench collects more than 200 tasks spanning linguistics, child development, mathematics, common-sense reasoning and social bias, and is explicitly designed to study both smooth and “breakthrough” capability gains as model scale increases Srivastava et al. (2022). Holistic Evaluation of Language Models (HELM) shifts the focus from single metrics to a scenario-based, multi-metric view, jointly reporting accuracy, calibration, robustness, fairness, bias, toxicity, and efficiency for a broad set of models Liang et al. (2022). TruthfulQA shows that larger models can be less truthful on adversarial questions that elicit human misconceptions Lin et al. (2022), while BBQ systematically probes stereotype-driven errors in question answering across several social dimensions Parrish et al. (2022). In line with these works, this study introduces TurkBench, a comprehensive Turkish-specific benchmark with 8,151 data samples across 21 subtasks.
Multilingual Benchmarks
Beyond English, several multilingual and low-resource benchmarks examine cross-lingual generalization. XTREME evaluates multilingual encoders across 40 typologically diverse languages and nine tasks, and shows a substantial gap between English and many lower-resourced languages, particularly for syntactic and retrieval tasks Hu et al. (2020). XGLUE defines 11 cross-lingual tasks in 19 languages, enforcing zero-shot transfer by providing training data only in English and evaluating on multiple target languages Liang et al. (2020). These multilingual benchmarks typically include Turkish as one language among many and often rely on translated or repurposed resources. By contrast, TurkBench is constructed entirely from Turkish sources, using expert-curated material from national exams, university coursework, sociology texts, Turkish literature, and real-world platforms. We explicitly avoid adapting existing NLP datasets or relying on synthetic data, which allows it to better capture language- and culture-specific phenomena in Turkish.
Benchmarks for Turkish
Mukayese is one of the first benchmarks for Turkish, which assembles datasets for tasks such as text classification, named entity recognition, and sentiment analysis Safaya et al. (2022). TurkishMMLU Yüksel et al. (2024) and TR-MMLU Bayram et al. (2025) adapt the MMLU paradigm to Turkish language, providing large-scale multiple-choice question answering benchmarks that reflect national educational standards; yet these resources primarily focus on exam-style multiple-choice question tasks. TR-MTEB focuses on sentence embeddings and introduces the first large-scale benchmark for Turkish sentence representations, covering six task families over 26 datasets Baysan et al. (2025). Turkish-PLU (Procedural Language Understanding) constructs a corpus of Turkish WikiHow texts, and defines tasks such as action linking and summarization; showing that Turkish-specific models outperform multilingual baselines on procedural understanding tasks Uzunoğlu and Şahin (2023). Cetvel is a recent benchmark covering 23 tasks grouped into seven categories based on existing publicly available datasets Er et al. (2025). Compared to these efforts, TurkBench differs in several key aspects. (i) All 8,151 instances across 21 subtasks are newly curated from high-quality Turkish sources with strategic partnerships with national institutions and university departments. (ii) All data is validated by human experts for correctness, grammatical well-formedness, and cultural sensitivity. (iii) We cover not only standard language understanding and classification but also reasoning, safety-oriented content moderation, and instruction following, which provides a much broader view of Turkish LLM capabilities. (iv) TurkBench is integrated into a public leaderboard that supports near-automatic submission and standardized evaluation.
Linguistic Perspective
There is increasing interest in benchmarks that explicitly target grammatical and cognitive phenomena. BLiMP (Benchmark of Linguistic Minimal Pairs) introduces 67 minimal-pair datasets that isolate contrasts in syntax, morphology and semantics, and evaluates whether LLMs prefer grammatical over ungrammatical sentences Warstadt et al. (2020). SyntaxGym builds on psycholinguistic experimental design and provides standardized suites for targeted syntactic evaluation Gauthier et al. (2020). CRAFT extends this idea to multimodal causal reasoning, introducing a video question-answering benchmark that tests understanding of interactions and counterfactuals Ates et al. (2022). For Turkish, surveys of available corpora and lexical resources emphasize important gaps in areas that are typologically salient, such as agglutinative morphology and relatively free word order Çöltekin et al. (2023). TurkBench is aligned with this linguistically informed line of work. Its tasks include rare and loanwords, idioms and metaphors, and part-of-speech classification; while other categories probe semantic textual similarity, natural language inference, and culturally grounded general knowledge. These tasks capture important structures and culturally specific content, while still being framed as realistic LLM tasks suitable for modern generative evaluation.
Table 1: The distribution of all tasks in TurkBench.
Task Instances
Turkish General Knowledge 200
MMLU 2,373
Reading Comprehension 482
Natural Language Inference 256
Summarization 262
Semantic Textual Similarity 225
Mathematical Reasoning 500
Complex Reasoning 100
Commonsense Reasoning 241
Sentiment Analysis 123
Topic Detection 240
Toxicity Detection 250
Bias Detection 250
Hallucination: Truthfulness 250
Hallucination: Faithfulness 250
Turkish Vocabulary: Rare Words 139
Turkish Vocabulary: Loan Words 165
Named Entity Recognition 438
Part-of-Speech 260
Metaphors and Idioms 150
Instruction Following 997
Total (21 tasks) 8,151
3 Benchmark Tasks
We explain each task briefly in this section. The summary statistics are listed in Table 1. The details of data construction, prompts, and data samples for all tasks are given in Appendix 9.2.
3.1 Knowledge
Turkish General Knowledge
This tasks reflects the culture, history, and daily life of Turkish people in general. It is expected that some words used here may be unique to Turkish people. It aims to test the general knowledge that is unique and relevant to Turkish people. There are 200 multiple-choice questions and answers, manually authored by domain experts. The evaluation metric is accuracy.
MMLU
The task is designed to assess general world knowledge and cross-disciplinary reasoning of large language models, inspired by the Massive Multitask Language Understanding benchmark (Hendrycks et al., 2020). There are multiple choice questions in a total of 24 distinct subject areas, chosen to reflect both secondary and tertiary education as well as key professional domains. 2,373 multiple choice questions are curated exclusively by having permissions from OSYM (The Turkish Measuring, Selection and Placement Center)-administered national exams and a complementary set of midterm/final assessments at Middle East Technical University. The evaluation metric is accuracy.
3.2 Language Understanding
Reading Comprehension
The task aims to assess a model’s ability to extract meaning from a text. In this task, the model is expected to identify essential information within the provided texts and answer questions. There are 482 open-ended questions curated by the domain experts from Hacettepe University, Sociology Department using the Sociology text sources such as academic papers, books, and other documents (will be referred to as "the Sociology experts" in the following sections) [^2] The methodology and workflow of the Sociology experts for this benchmark is explained at Appendix 9.3.. The evaluation metric is LLM-as-a-Judge.
Natural Language Inference
The task measures a model’s ability to understand the relationships between two sentences. In this task, the model is asked to determine the valid relationship between a given premise and hypothesis sentence. This relationship is classified into one of three categories: entailment (true), contradiction (false), or neutral (uncertain). There are 256 multiple-choice questions curated by the domain experts using the Sociology experts. The evaluation metric is accuracy.
Summarization
The task measures a model’s ability to condense a text while presenting its main ideas and most important information. In this task, the model is asked to read a long Turkish text and summarize its key elements. The model is expected to create a concise and clear summary that retains the essential information of the text while omitting unnecessary details. In this task, the model is expected to perform summarization in accordance with Turkish sentence structure and present the main idea of the text in a brief and clear manner without losing its essence. There are 262 open-ended questions curated by the Sociology experts. The evaluation metric is LLM-as-a-Judge.
Semantic Textual Similarity
The task aims to evaluate the semantic similarity between two sentences. In this task, the model is expected to assess the degree of similarity between two Turkish sentences on a scale between 1 and 5. The task determines whether the sentences convey the exact same meaning, are partially similar, or are entirely different. The model should accurately measure the semantic closeness between the two sentences, taking into account idiomatic expressions and indirect expressions commonly used in Turkish. In doing so, the model should go beyond surface-level similarity and consider nuanced differences between the sentences. There are 225 single-score questions curated by the Sociology experts. The evaluation metric is Pearson Correlation Coefficient.
3.3 Reasoning
Mathematical Reasoning
The mathematics task aims to evaluate a model’s ability in numerical reasoning, problem-solving, and understanding of mathematical concepts. In this task, the model is expected to correctly solve questions covering topics such as basic arithmetic, algebra, geometry, and probability. The mathematics task in Turkish tests the model’s capability to comprehend both the mathematical content and its presentation in the Turkish language. In this task, the model is expected to solve Turkish mathematical questions accurately, making both numerical and logical inferences. Turkish mathematics tasks reveal the model’s abilities in numerical reasoning and logical thinking as well as its linguistic understanding. There are 500 single-score questions based on the questions extracted from the TUBITAK (The Scientific and Technological Research Council of Turkey) Science Olympiad, and Middle East Technical University exams. The evaluation metric is accuracy.
Complex Reasoning
Complex tasks measure a model’s ability to use multiple skills simultaneously to solve multifaceted and challenging problems. These tasks typically require language comprehension, logical reasoning, long text analysis, information synthesis, and the ability to process various data types. In Turkish complex tasks, the model is expected to solve long or multi-step problems accurately and derive conclusions by understanding the nuances and details within the content. This task evaluates the model’s capacity to go beyond surface-level understanding and perform deep analysis and comprehension. There are 100 multiple-choice questions based on the questions extracted from the OSYM (The Turkish Measuring, Selection and Placement Center)-administered national ALES exams. The evaluation metric is accuracy.
Commonsense Reasoning
The commonsense task aims to evaluate a model’s ability to make logical and expected inferences in daily life. In this task, the model is expected to interpret a situation or sentence based on real-world knowledge and logic. For Turkish commonsense tasks, it is essential that the model can respond using common cultural knowledge, widely accepted societal beliefs, and logical inferences within Turkish texts or dialogues. This task assesses the model’s ability to provide accurate answers based not only on language skills but also on general knowledge, life experience, and commonly accepted information. There are 241 multiple-choice questions curated by the Sociology experts. The evaluation metric is accuracy.
3.4 Content Moderation
Sentiment Analysis
The task aims to evaluate a model’s ability to determine the emotional tone expressed in texts. In this task, the model is expected to analyze a given Turkish text and identify whether it conveys a positive, negative, or neutral sentiment. This task requires the model to understand sentiment orientation within texts and correctly classify emotionally charged expressions. There are 123 multiple-choice questions collected from a diverse set of platforms, including Airbnb, X (formerly Twitter), Trendyol, Hepsiburada, Reddit, and YouTube. The evaluation metric is accuracy.
Topic Detection
The task aims to evaluate a model’s ability to identify a text’s main subject or theme. The model is expected to analyze a given Turkish text and determine the general topic or theme it belongs to. The task tests the model’s ability to recognize various topics in the Turkish language, comprehend the main idea in the text, and assign it to a meaningful category. This task demonstrates the model’s capacity to perform accurate classification beyond mere grammar. There are 240 multiple-choice questions crafted from the Sociology sources. The evaluation metric is accuracy.
Toxicity Detection
The task involves assessing the model’s outputs for harmful, offensive, or inappropriate content across several key categories. This evaluation process is crucial for ensuring the responsible development and deployment of AI systems. There are 250 open-ended questions curated by the Sociology experts. The evaluation metric is LLM-as-a-Judge.
Bias Detection
The task aims to evaluate a model’s ability to recognize and mitigate any biases present within its responses. In this task, the model is tested to identify and address any potential stereotypes, prejudices, or unjust assumptions that may be present in Turkish language data. The goal is to ensure that the model’s outputs are fair, balanced, and free of language or content that could perpetuate harmful stereotypes or reinforce societal biases. The Turkish bias detection task focuses on ensuring the model’s responses remain impartial and culturally sensitive, maintaining inclusivity across various demographic, social, and cultural dimensions. There are 250 multiple-choice questions curated by the Sociology experts. The evaluation metric is accuracy.
Hallucination: Truthfulness and Faithfulness
We include two tasks to evaluate hallucinations: Truthfulness and Faithfulness. The Truthfulness task aims to evaluate the LLM’s ability to access and utilize its internal knowledge base to answer factual questions correctly. It specifically focuses on areas where common misconceptions or "folk wisdom" might mislead. This helps assess if the LLM can distinguish between reliable information and popular beliefs that may be inaccurate. The Faithfulness task focuses on evaluating the LLM’s ability to comprehend and utilize the information presented within a given context. It is specifically useful to see LLM’s capabilities and potential for Retrieval Augmented Generation (RAG) Systems where the augmented context is given to the LLM. There are 250 open-ended questions for each (Truthfulness and Faithfulness), curated by the Sociology experts. The evaluation metric is LLM-as-a-Judge.
3.5 Turkish Grammar and Vocabulary
Turkish Vocabulary: Rare Words
The task aims to evaluate a model’s knowledge of Turkish vocabulary. The model is asked to find the synonym of the given word. The subset gives a rare word that is outside of the scope of the basic core vocabulary, and asks for its synonym. This way the model’s knowledge of Turkish vocabulary outside of the daily usage can be evaluated. There are 139 multiple-choice questions curated by the domain experts from Hacettepe University, Sociology Department using the Turkish Literature sources such as novels (will be referred to as "the Turkish Literature experts" in the following sections). The evaluation metric is accuracy.
Turkish Vocabulary: Loan Words
The loan words subset gives a word of foreign origin and asks for its synonym with Turkish origin. This way the model’s knowledge of the words with foreign and Turkish origin of the words can be evaluated. There are 165 multiple-choice questions curated by the Turkish Literature experts using various text sources such as the "Dil Derneği" articles and the “Türkçesi Varken” pamphlets. The evaluation metric is accuracy.
Token Classification: Named Entity Recognition
The task aims to measure a model’s ability to identify and classify specific entities within a text. The model is expected to accurately identify entities such as names of people, places, organizations, dates, and numbers in a given Turkish text and classify them into the correct categories. The task tests the model’s ability to correctly recognize and interpret proper nouns, locations, and other named entities in Turkish sentences. This task demonstrates the model’s capability to distinguish and correctly classify specific information beyond basic language processing. There are 438 multiple-choice questions curated by the authors using recent news articles. The evaluation metric is accuracy.
Token Classification: Part-of-Speech
The task aims to evaluate a model’s ability to categorize words in a text into grammatical categories. The model is expected to classify each word in a given Turkish text with the correct Part-of-Speech (POS) label, such as noun, verb, adjective, or adverb. The Turkish POS tagging task tests the model’s ability to accurately identify the grammatical function of words within Turkish sentence structures. This task demonstrates the model’s capacity to understand not only individual words but also sentence structure and grammatical rules. There are 260 multiple-choice questions curated by the authors using recent news articles. The evaluation metric is accuracy.
Metaphors and Idioms
The metaphors and idioms task aims to evaluate a model’s knowledge and understanding of Turkish metaphors and idioms. In this task we give a context to the model and ask it to fill or find the suiting idiom correctly. There are 150 multiple-choice questions curated by the Turkish Literature experts using the Turkish Idioms and Proverbs Dictionary. The evaluation metric is accuracy.
3.6 Instruction Following
The Instruction Following task aims to assess a model’s ability to understand, interpret, and accurately follow user instructions. The model is provided with specific instructions in Turkish and is expected to respond by fulfilling the instructions exactly as requested, without deviation. The task tests the model’s ability to accurately comprehend the intent, requirements, and nuances of the instructions, responding in a precise and contextually appropriate manner. It also assesses the model’s ability to handle different levels of complexity in instructions, from simple requests to more detailed, multi-step tasks. There are 997 open-ended questions curated by the partners at Bogazici University, Computer Engineering Department. The evaluation metric is LLM-as-a-Judge.
4 Experiments
In this section, we describe our experimental setup and evaluation metrics, along with the experimental results. The experiments are supported by Turkcell AI.
4.1 Experimental Design
Our benchmark covers a total of 21 tasks, consisting of 7 open-ended and 14 multiple-choice tasks. For the open-ended tasks, we adopt an LLM-as-a-Judge paradigm. In particular, we employ GPT-4o-mini as the judge model, which evaluates system responses based on semantic alignment with reference answers. This approach enables us to assess inherently subjective tasks (e.g., summarization, faithfulness, instruction following) that cannot be reduced to a single ground truth label. For the multiple-choice tasks, in contrast, objective evaluation is performed by directly comparing model predictions against gold answers.
Metaprompting and Prompt Selection
A key aspect of our design is the use of metaprompting for robust prompt optimization. Instead of manually crafting prompts, we provide the model with representative samples and ask it to generate candidate prompts tailored for each task. This approach is chosen because of the fact that it reduces bias stemming from human-designed prompts, allows systematic exploration of diverse phrasing strategies, and ensures fairness across tasks and models by relying on a data-driven procedure.
We conduct systematic prompt selection experiments by testing multiple candidates across different LLMs. Table 2 shows an example for Topic Detection, where six different prompt formulations are compared on the Gemma-3-1B model. Turkish-language prompts consistently outperform the English variant, highlighting the importance of designing prompts in the target evaluation language.
Table 2: Prompt selection results for the Topic Detection task on Gemma-3-1B. The best-performing prompt is highlighted.
Prompt Variant Accuracy (%)
U1* 25.10
U2 26.74
U3 24.27
U4 26.33
U5 (Best) 29.21
Ueng (English) 20.98
Systematic metaprompting produces significant gains across tasks. For instance, Turkish General Knowledge gets improved by over +13 points, while Sentiment Analysis and Topic Detection showed gains of +2–3 points after optimized prompts are selected. These improvements confirm that prompt phrasing significantly affects evaluation reliability in Turkish, and that data-driven prompt discovery is important. We also observe that Turkish-language prompts consistently outperform their English equivalents, underscoring the importance of native-language alignment.
Reproducibility
To guarantee reproducibility, all experiments are conducted under standardized conditions across models. This includes consistent random seeds, identical evaluation scripts, and controlled inference settings (e.g., temperature, maximum token limits). Together, these design choices yield a transparent experimental framework that accounts for prompt sensitivity, data difficulty, and evaluation fairness across the entire benchmark.
4.2 Evaluation Metrics
Evaluation in TurkBench follows a task-dependent methodology. For multiple-choice tasks (e.g., Turkish General Knowledge, MMLU, Topic Detection), we adopt exact match accuracy as the evaluation metric, where a prediction is counted as correct only if the output exactly matches the gold-standard label (A, B, C, or D).
For open-ended tasks such as summarization, faithfulness, and instruction following, we employ an LLM-as-a-Judge strategy. Specifically, GPT-4o-mini is used as the evaluator, comparing system responses to references and providing judgments of semantic alignment, factual consistency, and instruction adherence. This enables reliable assessment of tasks where discrete gold labels are insufficient. This combined design ensures both objectivity (via exact match) and subjectivity-aware evaluation (via LLM-as-a-Judge).
In addition, we employ specialized metrics for some tasks. Faithfulness is measured with Deepeval’s Faithfulness Metric, which scores the consistency of generated answers with the given context. Toxicity is assessed using Deepeval’s Toxicity Metric, which evaluates harmfulness, bias, and cultural sensitivity. Bias Detection relies on accuracy with respect to curated ambiguous vs. disambiguated contexts, following the BBQ framework.
4.3 Experimental Results
Prompt Tuning
Prompt tuning is a technique used to adapt large language models for specific tasks without retraining. We manually refine the prompts and observe the performance improvements. The effectiveness of prompt tuning is evidenced by quantitative gains across tasks. Table 3 summarizes improvements after prompt refinement, where we observe consistent accuracy increments, including +13.71 points in Turkish General Knowledge and +3.48 points in MMLU.
Table 3: Prompt tuning improvements across selected tasks.
Task Old New Imp.
Sentiment Analysis 18.72 21.19 +2.37
Turkish Gen. Knowl. 10.03 23.74 +13.71
Topic Detection 15.36 18.65 +3.29
MMLU 13.86 17.34 +3.48
Complex Reasoning 17.62 20.83 +3.21
Table 4: Accuracy by difficulty levels on Turkish General Knowledge.
Model Full w/o 1 w/o 1&2 only 1&2
Gemma-2-9B 62.33 65.51 78.94 60.18
Gemma-3-27B 74.65 72.41 81.57 73.14
Qwen-3-0.6B 23.28 25.28 31.58 19.44
Qwen-3-1.7B 27.39 22.98 26.32 25.92
Table 5: TurkBench Results: Comprehensive Evaluation Across All Tasks
Model TK ML RC NL SM ST MA CX CS BI TR FA TX IF SA TD VO NE PS ID Avg
gpt-oss-120b 77.6 72.2 96.1 80.0 81.8 90.4 26.6 80.8 94.7 82.8 75.3 85.1 99.9 93.6 48.9 85.6 85 70.0 78.2 68.0 78.6
GLM-4.6 83.9 76.0 94.0 85.5 76.5 88.2 34.4 48.1 63.1 88.6 67.0 87.8 95.6 92.6 49.6 84.4 96 69.2 79.0 78.0 76.9
DeepSeek-V3.1 77.6 46.7 95.1 28.2 80.1 91.5 58.6 66.3 91.8 79.6 77.2 87.4 98.7 94.9 48.9 84.8 93 56.5 83.6 63.3 75.2
Qwen3-Next-80B-Inst 67.2 61.4 94.4 85.1 81.7 93.0 63.4 26.9 92.2 74.2 67.9 88.7 95.3 94.3 46.7 85.6 80 58.3 81.3 61.3 75.0
Qwen3-30B-A3B-Inst 65.1 58.9 90.8 81.6 81.9 86.9 55.4 22.1 90.6 78.2 62.9 83.4 97.5 92.5 44.5 95.2 89 55.0 78.1 68.7 73.4
gemma-3-27b-it 74.7 53.5 93.0 75.7 74.1 90.2 19.6 27.9 91.4 80.1 61.8 84.4 94.0 84.8 42.6 88.1 90 66.0 84.4 85.3 73.0
Qwen3-235B-Instruct 73.4 61.6 95.4 33.7 81.8 92.2 47.2 24.0 91.4 73.2 77.4 89.1 98.1 94.5 47.4 87.7 77 58.5 81.0 59.3 72.2
gemma-3-12b-TR-V1 71.4 48.0 92.6 71.4 77.0 92.6 23.4 26.9 91.0 77.8 60.1 86.2 93.0 83.7 46.7 80.3 90 63.9 73.4 74.7 71.2
gemma-3-12b-it 71.4 48.2 92.6 71.4 77.5 92.6 22.4 23.1 91.4 78.0 59.6 88.1 92.7 81.2 47.5 79.8 91 63.5 73.6 74.7 71.0
Qwen3-235B-A22B 66.1 61.3 94.0 29.4 71.5 93.3 33.8 15.4 88.5 76.0 60.0 86.1 98.7 91.2 43.1 86.4 57 55.8 77.7 79.3 68.2
Qwen2.5-14B-Instruct 58.9 45.2 89.0 82.4 75.0 91.9 18.8 27.9 90.2 82.8 51.1 83.4 99.0 64.5 46.0 68.7 80 51.6 64.3 60.0 66.5
Tongyi-DR-30B 64.6 35.3 85.6 75.7 72.9 88.0 29.0 1.0 85.7 84.2 59.5 85.6 98.3 59.9 46.0 76.1 83 55.1 54.0 68.0 65.4
TR-Gemma-9b-v0.1 67.7 48.0 92.5 62.0 78.3 89.1 9.8 23.1 89.3 79.2 65.1 88.3 96.7 83.3 46.0 80.3 84 0.0 46.8 77.3 65.3
Qwen3-32B 66.7 38.0 93.0 81.6 72.8 90.9 0.0 26.0 91.0 8.6 61.0 87.6 98.0 83.8 48.9 85.2 71 60.4 77.7 50.7 64.6
gemma-2-9b-it 62.3 39.4 90.3 76.5 74.4 86.8 6.4 8.7 89.3 18.8 54.2 86.0 96.1 73.6 39.5 84.8 81 48.8 69.1 68.0 62.7
aya-expanse-8b 55.7 32.5 90.0 40.8 77.0 81.1 4.2 23.1 86.9 24.2 54.8 87.0 98.4 82.2 41.6 71.6 78 38.6 63.9 40.0 58.6
Qwen2.5-7B-Instruct 39.1 29.4 85.3 53.7 73.4 91.8 10.8 23.1 74.6 68.2 43.4 82.4 97.9 61.8 23.4 70.0 39 34.8 63.8 32.7 54.9
Llama-3.1-8B-Instruct 40.1 19.0 89.5 31.0 71.5 81.9 2.8 25.0 33.2 36.8 45.3 79.2 98.6 71.9 21.2 28.4 38 26.7 61.6 13.3 45.7
DeepSeek-R1-Q3-8B 31.3 20.9 84.0 32.2 68.3 54.1 0.2 21.2 50.8 12.0 55.7 80.0 98.5 73.1 22.6 39.5 24 44.5 45.8 22.0 44.0
Qwen3-14B 35.4 29.5 72.2 32.2 68.3 61.2 14.6 5.8 80.7 44.0 21.1 66.7 98.0 32.1 30.7 58.8 44 19.7 18.0 26.7 43.0
Phi-4-mini-instruct 3.7 10.3 84.9 60.0 65.4 89.3 6.2 15.4 17.6 12.2 31.8 74.9 98.4 62.1 37.2 46.9 41 32.3 31.3 21.3 42.1
gemma-2-2b-it 24.5 6.4 86.1 32.2 70.2 84.2 2.0 18.3 76.2 33.2 37.6 83.4 96.6 59.1 0.0 0.4 0 31.9 29.4 1.3 38.6
Magistral-Small-2506 49.5 27.4 57.8 3.9 53.0 81.7 11.4 7.7 38.1 25.3 16.6 82.6 96.1 10.7 21.9 68.7 15 27.4 57.7 12.7 38.3
Qwen3-1.7B 27.4 21.3 52.5 41.2 62.6 51.0 0.1 20.2 55.3 10.9 22.3 82.7 98.1 10.1 30.9 17.3 23 36.6 40.0 26.0 36.5
TDM-8b-v0.1 23.4 21.3 51.3 34.1 66.8 0.0 2.0 18.3 58.2 0.0 25.2 87.8 95.7 5.0 24.1 29 0.0 0.0 27.3 30.0
Qwen3-0.6B 23.3 19.1 35.5 30.2 32.4 49.7 0.0 18.3 65.6 8.9 12.6 69.2 100.0 4.5 19.1 10.7 26 20.9 17.2 26.0 29.5
Kumru-2B 2.7 10.6 79.0 1.6 54.7 58.0 0.2 0.0 8.6 29.2 49.3 76.5 97.4 47.0 11.0 6.2 12 0.0 0.4 2.0 27.3
Task Abbreviations: TK = Turkish General Knowledge, ML = MMLU Topics, RC = Reading Comprehension, NL = Natural Language Inference, SM = Summarization, ST = Semantic Textual Similarity, MA = Mathematics, CX = Complex Reasoning, CS = Commonsense Reasoning, BI = Bias (MC), TR = Truthfulness, FA = Faithfulness, TX = Toxicity, IF = Instruction Following, SA = Sentiment Analysis, TD = Topic Detection, VO = Turkish Vocabulary, NE = Named Entity Recognition, PS = Parts of Speech, ID = Metaphors & Idioms, Avg = Overall Average. Models are sorted by average score (descending).
Difficulty Calibration
Another essential element of our design is the calibration of data difficulty levels. Many benchmark tasks, such as Turkish General Knowledge, contain questions labeled with difficulty scores. To ensure balanced evaluation, we experiment with including and excluding certain difficulty ranges. Table 4 illustrates this process by showing how model accuracy shifts when easy questions (levels 1 and 2) are removed.
Difficulty-level filtering experiments reveal that benchmark reliability strongly depends on balanced sampling. When easy questions (levels 1 and 2) are removed, models such as Gemma-3-27B display higher relative accuracy, while smaller models lost performance more significantly. This indicates that evaluation without difficulty balancing risks overestimating model competence.
Comparative Results
We evaluate and compare 27 open-source large language model in the TurkBench benchmark, sorted by the average score across all tasks, in Table 5. The model details are given in Appendix 9.4.
For tasks such as summarization and faithfulness, LLM-as-a-Judge performs well. Ground-truth references alone can not capture nuances of coherence and factuality. Judge-based evaluation enables meaningful differentiation between models, and also reveals weaknesses in smaller models that exact-match metrics would have missed.
Larger models (e.g., Qwen-32B, Gemma-27B) consistently outperform smaller ones across both multiple-choice and open-ended tasks. However, even state-of-the-art systems struggle with culturally grounded reasoning (idioms, proverbs, Turkish-specific general knowledge), where accuracy remains well below English-centric benchmarks.
5 Leaderboard
We release TurkBench’s public leaderboard [^3] https://huggingface.co/turkbench to publish results in a transparent and easily navigable manner. Accordingly, we adopt three guiding principles: (i) Reliance on an open-source technology stack, (ii) near-automatic model-submission and evaluation, and (iii) a lightweight and intuitive user interface. The details of the leaderboard implementation are given in Appendix 9.5.
6 Conclusion
This study introduces TurkBench, a comprehensive and culturally-aware benchmark designed to evaluate the capabilities of large language models in the Turkish language. TurkBench addresses a critical gap in Turkish LLM evaluation, which is mostly known as English-centric. TurkBench offers researchers and developers a robust tool to assess the true performance of models on tasks that capture the unique agglutinative morphology and syntactic rules of Turkish, moving beyond simple translations of existing English benchmarks.
TurkBench not only facilitates more accurate comparisons between different Turkish large language models, but also promotes the development of more responsible and reliable AI systems for Turkish-speaking users. The public release of its leaderboard aims to foster a competitive environment by accelerating progress within the Turkish AI ecosystem, and guiding future advancements in language model evaluation for low-resource languages. Future work would include the development of similar benchmarks in the ethical evaluation of large language models in low-resource settings. Domain-specific benchmarks are another opportunity for further development.
7 Limitations
While TurkBench represents an important advancement in evaluating Turkish large language models, there exist some limitations. The benchmark’s data, while high-quality, is mainly sourced from formal, academic, and journalistic contexts such as national exams, university materials, and established news outlets. This focus on a standardized register of Turkish means that the benchmark may not fully capture a model’s proficiency in handling informal language, regional dialects, or the dynamic slang observed on social media.
Another limitation stems from the evaluation methodology for open-ended tasks. The study employs the LLM-as-a-Judge approach, using GPT-4o-mini to assess subjective tasks like summarization and instruction following. While this method allows for scalable and consistent evaluation, it is not without potential drawbacks. The judge model may have its own inherent biases, stylistic preferences, or gaps in understanding deep cultural nuances specific to Turkish, which could influence its evaluations.
Furthermore, the scope of TurkBench, while broad with 21 distinct subtasks, is confined to text-based evaluation. It does not currently assess multimodal capabilities, such as understanding images and text together, or speech-based tasks like automatic speech recognition and text-to-speech in Turkish.
8 Ethical Considerations
A primary ethical consideration in the development of TurkBench is the sourcing of data. The benchmark is constructed from a wide array of sources, including national examinations, academic materials, news articles, and user-generated content from public platforms such as Reddit and YouTube. While many of these sources are public, they can contain personally identifiable information or opinions of individuals who did not explicitly consent to their data being used in a benchmark. We have an ethical responsibility to ensure that the data is handled in a way that minimizes privacy risks, such as by anonymizing data where possible and avoiding the inclusion of sensitive personal details that are not essential for the evaluation task.
By designing tasks to identify stereotypes, prejudices, and offensive content, the benchmark encourages developers to build models that are not only linguistically proficient but also fair and culturally sensitive. The methodology for bias detection provides a structured way to quantify and address a critical failure mode in language models. This focus on safety and responsibility is crucial for developing AI systems that serve Turkish-speaking users reliably.
The creation of a comprehensive benchmark such as TurkBench also introduces considerations regarding its potential for misuse. While the intended purpose is to foster improvement and safety, the benchmark could be used by malicious actors to fine-tune and perfect models for harmful applications, such as generating highly convincing misinformation, propaganda, or social engineering attacks that are culturally and linguistically tailored to a Turkish audience. Furthermore, there is a risk that the specific definitions of "bias" and "toxicity" embedded within the benchmark, though curated by experts, may not encompass all forms of harm and could reflect the specific perspectives of its creators.
Generative AI is used in writing of this study to assist with language editing. All scientific contributions, data construction, data analysis, and interpretations presented in this work are original and were conducted entirely by the authors.
Acknowledgments:
We would like to thank to OSYM (The Turkish Measuring, Selection and Placement Center), TUBITAK (The Scientific and Technological Research Council of Turkey), and Turkcell AI for their permissions to use certain resources in this benchmark.

### v2

@@
 Turkish, with its agglutinative morphology and distinct syntactic rules, presents unique challenges that cannot be easily assessed by frameworks designed for Indo-European languages Hakkani-Tür et al.
-(2002); Oflazer (2014).
+(2002); Oflazer (2014); Toraman (2024).
 Although there are efforts to develop benchmarks for Turkish LLMs Safaya et al.
@@
 (2024); Er et al.
-(2025), these resources primarily focus on exam-style tasks or do not cover many capabilities that drive real-world use of Turkish LLMs, such as open-ended instruction following, safety and content moderation, or fine-grained grammar and vocabulary control.
-Moreover, such benchmarks are mostly designed by using existing publicly available datasets, and lack of public leaderboards serving automated evaluation.
+(2025), these resources are mostly designed by using existing publicly available datasets, and lack of public leaderboards accepting online submissions for automated evaluation.
+Moreover, existing benchmarks do not cover a wide range of capabilities that drive real-world use of Turkish LLMs, such as open-ended instruction following, safety and content moderation, or fine-grained grammar and vocabulary control.
 In this study, we present TurkBench, which is a comprehensive benchmark specifically designed by curating novel data for the Turkish language.
@@
 Turkish-PLU (Procedural Language Understanding) constructs a corpus of Turkish WikiHow texts, and defines tasks such as action linking and summarization; showing that Turkish-specific models outperform multilingual baselines on procedural understanding tasks Uzunoğlu and Şahin (2023).
+There are also some efforts to develop the Turkish versions of some existing English benchmarks Acikgoz et al.
+(2024).
 Cetvel is a recent benchmark covering 23 tasks grouped into seven categories based on existing publicly available datasets Er et al.
@@
 For the multiple-choice tasks, in contrast, objective evaluation is performed by directly comparing model predictions against gold answers.
+We do not share the prompts used for LLM-as-a-Judge due to online leaderboard evaluation.
 Metaprompting and Prompt Selection
@@
 Table 5: TurkBench Results: Comprehensive Evaluation Across All Tasks
-Model TK ML RC NL SM ST MA CX CS BI TR FA TX IF SA TD VO NE PS ID Avg
-gpt-oss-120b 77.6 72.2 96.1 80.0 81.8 90.4 26.6 80.8 94.7 82.8 75.3 85.1 99.9 93.6 48.9 85.6 85 70.0 78.2 68.0 78.6
-GLM-4.6 83.9 76.0 94.0 85.5 76.5 88.2 34.4 48.1 63.1 88.6 67.0 87.8 95.6 92.6 49.6 84.4 96 69.2 79.0 78.0 76.9
-DeepSeek-V3.1 77.6 46.7 95.1 28.2 80.1 91.5 58.6 66.3 91.8 79.6 77.2 87.4 98.7 94.9 48.9 84.8 93 56.5 83.6 63.3 75.2
-Qwen3-Next-80B-Inst 67.2 61.4 94.4 85.1 81.7 93.0 63.4 26.9 92.2 74.2 67.9 88.7 95.3 94.3 46.7 85.6 80 58.3 81.3 61.3 75.0
-Qwen3-30B-A3B-Inst 65.1 58.9 90.8 81.6 81.9 86.9 55.4 22.1 90.6 78.2 62.9 83.4 97.5 92.5 44.5 95.2 89 55.0 78.1 68.7 73.4
-gemma-3-27b-it 74.7 53.5 93.0 75.7 74.1 90.2 19.6 27.9 91.4 80.1 61.8 84.4 94.0 84.8 42.6 88.1 90 66.0 84.4 85.3 73.0
-Qwen3-235B-Instruct 73.4 61.6 95.4 33.7 81.8 92.2 47.2 24.0 91.4 73.2 77.4 89.1 98.1 94.5 47.4 87.7 77 58.5 81.0 59.3 72.2
-gemma-3-12b-TR-V1 71.4 48.0 92.6 71.4 77.0 92.6 23.4 26.9 91.0 77.8 60.1 86.2 93.0 83.7 46.7 80.3 90 63.9 73.4 74.7 71.2
-gemma-3-12b-it 71.4 48.2 92.6 71.4 77.5 92.6 22.4 23.1 91.4 78.0 59.6 88.1 92.7 81.2 47.5 79.8 91 63.5 73.6 74.7 71.0
-Qwen3-235B-A22B 66.1 61.3 94.0 29.4 71.5 93.3 33.8 15.4 88.5 76.0 60.0 86.1 98.7 91.2 43.1 86.4 57 55.8 77.7 79.3 68.2
-Qwen2.5-14B-Instruct 58.9 45.2 89.0 82.4 75.0 91.9 18.8 27.9 90.2 82.8 51.1 83.4 99.0 64.5 46.0 68.7 80 51.6 64.3 60.0 66.5
-Tongyi-DR-30B 64.6 35.3 85.6 75.7 72.9 88.0 29.0 1.0 85.7 84.2 59.5 85.6 98.3 59.9 46.0 76.1 83 55.1 54.0 68.0 65.4
-TR-Gemma-9b-v0.1 67.7 48.0 92.5 62.0 78.3 89.1 9.8 23.1 89.3 79.2 65.1 88.3 96.7 83.3 46.0 80.3 84 0.0 46.8 77.3 65.3
-Qwen3-32B 66.7 38.0 93.0 81.6 72.8 90.9 0.0 26.0 91.0 8.6 61.0 87.6 98.0 83.8 48.9 85.2 71 60.4 77.7 50.7 64.6
-gemma-2-9b-it 62.3 39.4 90.3 76.5 74.4 86.8 6.4 8.7 89.3 18.8 54.2 86.0 96.1 73.6 39.5 84.8 81 48.8 69.1 68.0 62.7
-aya-expanse-8b 55.7 32.5 90.0 40.8 77.0 81.1 4.2 23.1 86.9 24.2 54.8 87.0 98.4 82.2 41.6 71.6 78 38.6 63.9 40.0 58.6
-Qwen2.5-7B-Instruct 39.1 29.4 85.3 53.7 73.4 91.8 10.8 23.1 74.6 68.2 43.4 82.4 97.9 61.8 23.4 70.0 39 34.8 63.8 32.7 54.9
-Llama-3.1-8B-Instruct 40.1 19.0 89.5 31.0 71.5 81.9 2.8 25.0 33.2 36.8 45.3 79.2 98.6 71.9 21.2 28.4 38 26.7 61.6 13.3 45.7
-DeepSeek-R1-Q3-8B 31.3 20.9 84.0 32.2 68.3 54.1 0.2 21.2 50.8 12.0 55.7 80.0 98.5 73.1 22.6 39.5 24 44.5 45.8 22.0 44.0
-Qwen3-14B 35.4 29.5 72.2 32.2 68.3 61.2 14.6 5.8 80.7 44.0 21.1 66.7 98.0 32.1 30.7 58.8 44 19.7 18.0 26.7 43.0
-Phi-4-mini-instruct 3.7 10.3 84.9 60.0 65.4 89.3 6.2 15.4 17.6 12.2 31.8 74.9 98.4 62.1 37.2 46.9 41 32.3 31.3 21.3 42.1
-gemma-2-2b-it 24.5 6.4 86.1 32.2 70.2 84.2 2.0 18.3 76.2 33.2 37.6 83.4 96.6 59.1 0.0 0.4 0 31.9 29.4 1.3 38.6
-Magistral-Small-2506 49.5 27.4 57.8 3.9 53.0 81.7 11.4 7.7 38.1 25.3 16.6 82.6 96.1 10.7 21.9 68.7 15 27.4 57.7 12.7 38.3
-Qwen3-1.7B 27.4 21.3 52.5 41.2 62.6 51.0 0.1 20.2 55.3 10.9 22.3 82.7 98.1 10.1 30.9 17.3 23 36.6 40.0 26.0 36.5
-TDM-8b-v0.1 23.4 21.3 51.3 34.1 66.8 0.0 2.0 18.3 58.2 0.0 25.2 87.8 95.7 5.0 24.1 29 0.0 0.0 27.3 30.0
-Qwen3-0.6B 23.3 19.1 35.5 30.2 32.4 49.7 0.0 18.3 65.6 8.9 12.6 69.2 100.0 4.5 19.1 10.7 26 20.9 17.2 26.0 29.5
-Kumru-2B 2.7 10.6 79.0 1.6 54.7 58.0 0.2 0.0 8.6 29.2 49.3 76.5 97.4 47.0 11.0 6.2 12 0.0 0.4 2.0 27.3
-Task Abbreviations: TK = Turkish General Knowledge, ML = MMLU Topics, RC = Reading Comprehension, NL = Natural Language Inference, SM = Summarization, ST = Semantic Textual Similarity, MA = Mathematics, CX = Complex Reasoning, CS = Commonsense Reasoning, BI = Bias (MC), TR = Truthfulness, FA = Faithfulness, TX = Toxicity, IF = Instruction Following, SA = Sentiment Analysis, TD = Topic Detection, VO = Turkish Vocabulary, NE = Named Entity Recognition, PS = Parts of Speech, ID = Metaphors & Idioms, Avg = Overall Average.
+Model Knowl. Lang. Understand. Reasoning Content Moderation Grammar IF Avg
+TK ML RC NL SM ST MA CX CS BI TR FA TX SA TD VO NE PS ID IF
+gpt-oss-120b 77.6 72.2 96.1 80.0 81.8 90.4 26.6 80.8 94.7 82.8 75.3 85.1 99.9 48.9 85.6 85 70.0 78.2 68.0 78.6 93.6
+GLM-4.6 83.9 76.0 94.0 85.5 76.5 88.2 34.4 48.1 63.1 88.6 67.0 87.8 95.6 49.6 84.4 96 69.2 79.0 78.0 76.9 92.6
+DeepSeek-V3.1 77.6 46.7 95.1 28.2 80.1 91.5 58.6 66.3 91.8 79.6 77.2 87.4 98.7 48.9 84.8 93 56.5 83.6 63.3 75.2 94.9
+Qwen3-80B-Inst 67.2 61.4 94.4 85.1 81.7 93.0 63.4 26.9 92.2 74.2 67.9 88.7 95.3 46.7 85.6 80 58.3 81.3 61.3 75.0 94.3
+Qwen3-30B-Inst 65.1 58.9 90.8 81.6 81.9 86.9 55.4 22.1 90.6 78.2 62.9 83.4 97.5 44.5 95.2 89 55.0 78.1 68.7 73.4 92.5
+gemma-3-27b-it 74.7 53.5 93.0 75.7 74.1 90.2 19.6 27.9 91.4 80.1 61.8 84.4 94.0 42.6 88.1 90 66.0 84.4 85.3 73.0 84.8
+Qwen3-235B-Inst 73.4 61.6 95.4 33.7 81.8 92.2 47.2 24.0 91.4 73.2 77.4 89.1 98.1 47.4 87.7 77 58.5 81.0 59.3 72.2 94.5
+gemma-3-12b-TR 71.4 48.0 92.6 71.4 77.0 92.6 23.4 26.9 91.0 77.8 60.1 86.2 93.0 46.7 80.3 90 63.9 73.4 74.7 71.2 83.7
+gemma-3-12b-it 71.4 48.2 92.6 71.4 77.5 92.6 22.4 23.1 91.4 78.0 59.6 88.1 92.7 47.5 79.8 91 63.5 73.6 74.7 71.0 81.2
+Qwen3-235B 66.1 61.3 94.0 29.4 71.5 93.3 33.8 15.4 88.5 76.0 60.0 86.1 98.7 43.1 86.4 57 55.8 77.7 79.3 68.2 91.2
+Qwen2.5-14B-Inst 58.9 45.2 89.0 82.4 75.0 91.9 18.8 27.9 90.2 82.8 51.1 83.4 99.0 46.0 68.7 80 51.6 64.3 60.0 66.5 64.5
+Tongyi-DR-30B 64.6 35.3 85.6 75.7 72.9 88.0 29.0 1.0 85.7 84.2 59.5 85.6 98.3 46.0 76.1 83 55.1 54.0 68.0 65.4 59.9
+TR-Gemma-9b 67.7 48.0 92.5 62.0 78.3 89.1 9.8 23.1 89.3 79.2 65.1 88.3 96.7 46.0 80.3 84 0.0 46.8 77.3 65.3 83.3
+Qwen3-32B 66.7 38.0 93.0 81.6 72.8 90.9 0.0 26.0 91.0 8.6 61.0 87.6 98.0 48.9 85.2 71 60.4 77.7 50.7 64.6 83.8
+gemma-2-9b-it 62.3 39.4 90.3 76.5 74.4 86.8 6.4 8.7 89.3 18.8 54.2 86.0 96.1 39.5 84.8 81 48.8 69.1 68.0 62.7 73.6
+aya-expanse-8b 55.7 32.5 90.0 40.8 77.0 81.1 4.2 23.1 86.9 24.2 54.8 87.0 98.4 41.6 71.6 78 38.6 63.9 40.0 58.6 82.2
+Qwen2.5-7B-Inst 39.1 29.4 85.3 53.7 73.4 91.8 10.8 23.1 74.6 68.2 43.4 82.4 97.9 23.4 70.0 39 34.8 63.8 32.7 54.9 61.8
+Llama-3.1-8B-Inst 40.1 19.0 89.5 31.0 71.5 81.9 2.8 25.0 33.2 36.8 45.3 79.2 98.6 21.2 28.4 38 26.7 61.6 13.3 45.7 71.9
+DeepSeek-Q3-8B 31.3 20.9 84.0 32.2 68.3 54.1 0.2 21.2 50.8 12.0 55.7 80.0 98.5 22.6 39.5 24 44.5 45.8 22.0 44.0 73.1
+Qwen3-14B 35.4 29.5 72.2 32.2 68.3 61.2 14.6 5.8 80.7 44.0 21.1 66.7 98.0 30.7 58.8 44 19.7 18.0 26.7 43.0 32.1
+Phi-4-mini-instruct 3.7 10.3 84.9 60.0 65.4 89.3 6.2 15.4 17.6 12.2 31.8 74.9 98.4 37.2 46.9 41 32.3 31.3 21.3 42.1 62.1
+gemma-2-2b-it 24.5 6.4 86.1 32.2 70.2 84.2 2.0 18.3 76.2 33.2 37.6 83.4 96.6 0.0 0.4 0 31.9 29.4 1.3 38.6 59.1
+Magistral-Small 49.5 27.4 57.8 3.9 53.0 81.7 11.4 7.7 38.1 25.3 16.6 82.6 96.1 21.9 68.7 15 27.4 57.7 12.7 38.3 10.7
+Qwen3-1.7B 27.4 21.3 52.5 41.2 62.6 51.0 0.1 20.2 55.3 10.9 22.3 82.7 98.1 30.9 17.3 23 36.6 40.0 26.0 36.5 10.1
+TDM-8b-v0.1 23.4 21.3 51.3 34.1 66.8 0.0 2.0 18.3 58.2 0.0 25.2 87.8 95.7 24.1 29 0.0 0.0 27.3 30.0 5.0
+Qwen3-0.6B 23.3 19.1 35.5 30.2 32.4 49.7 0.0 18.3 65.6 8.9 12.6 69.2 100 19.1 10.7 26 20.9 17.2 26.0 29.5 4.5
+Kumru-2B 2.7 10.6 79.0 1.6 54.7 58.0 0.2 0.0 8.6 29.2 49.3 76.5 97.4 11.0 6.2 12 0.0 0.4 2.0 27.3 47.0
+Task Abbreviations: TK = Turkish General Knowledge, ML = MMLU Topics, RC = Reading Comprehension, NL = Natural Language Inference, SM = Summarization, ST = Semantic Textual Similarity, MA = Mathematics, CX = Complex Reasoning, CS = Commonsense Reasoning, BI = Bias (MC), TR = Truthfulness, FA = Faithfulness, TX = Toxicity, SA = Sentiment Analysis, TD = Topic Detection, VO = Turkish Vocabulary, NE = Named Entity Recognition, PS = Parts of Speech, ID = Metaphors & Idioms, IF = Instruction Following, Avg = Overall Average.
 Models are sorted by average score (descending).
@@
 Comparative Results
-We evaluate and compare 27 open-source large language model in the TurkBench benchmark, sorted by the average score across all tasks, in Table 5. The model details are given in Appendix 9.4.
+We evaluate and compare 27 open-source large language model in the TurkBench benchmark, sorted by the average score across all tasks, in Table 5.
+The results are based on a single run of each model.
+The model details are given in Appendix 9.4.
 For tasks such as summarization and faithfulness, LLM-as-a-Judge performs well.
@@
 We would like to thank to OSYM (The Turkish Measuring, Selection and Placement Center), TUBITAK (The Scientific and Technological Research Council of Turkey), and Turkcell AI for their permissions to use certain resources in this benchmark.
+We would also like to thank to all data annotators worked in this project.